Solve the system S x1 | 5x1 +x2 +4x3 +6x2 +4x3 = = 7 2
Answer #1

Solve the system +x2 +4x3 5x +6х, +4хз 7 X1 2 = 니 L4 2. R2 - h2-sRI 20 33 -I6 = 4O-2 0X3 X2 = -33 + 16X3 free хз II rt rt